OpenStudy (anonymous):

If f(x) = 5x + 3, then what is f–1(x) ?

OpenStudy (precal):

 you are asking for the inverse of the function correct?

OpenStudy (anonymous):

yes, sorry about that.

OpenStudy (precal):

y=5x+3 rewrite using y for f(x) switch x and y x=5y+3 solve for y x=5y+3 x-3=5y (x-3)/5=y

OpenStudy (precal):

when you graph your original function with its inverse, the line of symmetry is y=x

OpenStudy (anonymous):

so you wouldn't have to factor anything?

OpenStudy (precal):

not here

OpenStudy (anonymous):

ok, thanks. that's probably why i kept messing up.

OpenStudy (precal):

just remember to switch your x and y variables and solve for y